OpenStudy (anonymous):

Solve the equation: 5x2 = −65

OpenStudy (anonymous):

are you sure you have typed it ryt?

OpenStudy (anonymous):

Is it 5x^2 or 5x(2)?

OpenStudy (anonymous):

5x^2=-65

OpenStudy (anonymous):

fistly take the 5 across and divide -65 by 5

OpenStudy (anonymous):

this leaves you with \[x^{2} = -13\]

OpenStudy (anonymous):

next you need to make \[x = \sqrt{-13}\] but this is not posible so we need to employ the imaginary number i therefore we have the answer as \[x =\sqrt{13}i\]

OpenStudy (anonymous):

ok thank you. :)

OpenStudy (anonymous):

no prob. Good luck.

OpenStudy (anonymous):

can you help w.another?